# Creation of Wavepackets for Quantum Chromodynamics on Quantum Computers

### v1

@@ III Interpolating Operators in One-dimensional Lattice Gauge Theories
We start by considering the simpler case of gauge theories in one space dimension, which is enough to show all the building blocks that are necessary in the physical case of quantum chromodynamics. The Kogut-Susskind Hamiltonians for lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$ in one space dimension can be compactly written as
@@ III.1 $U(1)$ Gauge Theory
In this model there are three stable mesons in the continuum limit, with associated interpolating operators
@@
In the staggered-fermion formulation these operators translate respectively to
$\Phi_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (20)
$\Phi_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (21)
$\Phi_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (22)
where we have dropped the factor $i$ in $\Phi_{3}$ as it is irrelevant for our purposes.
@@ III.2 $SU(2)$ Gauge Theory
Analogously to the $U(1)$ theory, the $SU(2)$ theory has mesons
$\Phi_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (34)
$\Phi_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (35)
$\Phi_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]}.$ (36)
@@
$\Phi_{4}(x)=\frac{1}{a}\epsilon_{\alpha\beta}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x),$ (37)
@@ III.3 $SU(3)$ Gauge Theory
The $SU(3)$ case is a more complicated version of the $SU(2)$ one. For the details see [35]. The basis states are labelled by eight quantum numbers
@@
$\Phi_{4}(x)=\frac{1}{a^{3/2}}\epsilon_{\alpha\beta\gamma}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x)\xi_{\gamma}^{\dagger}(x).$ (45)
@@ IV Interpolating Operators on Qubits
[…]
where $\sigma^{\pm}=(\sigma^{x}\pm i\sigma^{y})/2$.
On the other hand, the local Hilbert space of the gauge field needs to be truncated. For $U(1)$ this is quite simple: if we use $k$ qubits for a single link, then we impose the cut-off $\Lambda=2^{k-1}$ and we encode the states $|-\Lambda\rangle,|-\Lambda+1\rangle,\dots,|\Lambda-1\rangle$ in the computational basis. The operator $U$ needs to be adapted to the truncation. A common choice is to take
$U|\mathcal{E}\rangle=|\mathcal{E}-1\rangle,\qquad\text{for }\mathcal{E}=-\Lambda+1,\dots,\Lambda-1,$ (49)
@@
In this way $U$ is a permutation operator and can be implemented on a qubit system using $O(k^{3})$ gates and no ancillary qubit, or using $O(k^{2})$ gates and $k-1$ ancillary qubits. For $SU(2)$ and $SU(3)$ the situation is more complicated but similar in essence. We refer the reader to [35] for the details. After imposing a cut-off $\Lambda$, the permutation operators $M_{\alpha}^{\text{L}/\text{R}},\,J^{\pm}$ of $SU(2)$, and $M_{\alpha}^{\text{L}/\text{R}},\,P^{\pm},\,Q^{\pm},\,T_{\text{L}/\text{R}}^{\pm}$ of $SU(3)$, take the same form as $U$ in equation (49).
The diagonal operators $c_{\alpha\beta}^{\pm}$ in (29) and the ones in (41), like for instance $\hat{C}_{11}^{\text{L}(a)}\hat{C}_{11}^{\text{R}(a)}\hat{N}_{1}$, are not unitary. However, their entries lie between $-1$ and $1$, and are efficiently computable classically. We can implement them in the following way: suppose that the function $f(x)$, where $x$ can take $M$ values encoded in the computational basis, is efficiently computable with classical techniques, and take an ancillary register of $\eta=O(\log_{2}M)$ qubits initialized in the state
@@
We can now write the interpolating operators in the qubit system. For $U(1)$ they take the form
$a\Phi_{1}(x)$ $=\sigma^{+}(x)\sigma^{-}(x)-\sigma^{+}(x+1)\sigma^{-}(x+1),$ (53)
$a\Phi_{2}(x)$ $=\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)+\sigma^{-}(x+1)U(x)\sigma^{+}(x),$ (54)
$a\Phi_{3}(x)$ $=\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)-\sigma^{-}(x+1)U(x)\sigma^{+}(x).$ (55)
@@
$a\Phi_{1}(x)$ $=\sum_{\alpha}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1),$ (56)
$a\Phi_{2}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (57)
$a\Phi_{3}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (58)
$a\Phi_{4}(x)$ $=2\sigma_{1}^{-}(x)\sigma_{2}^{-}(x),$ (59)
@@
$\zeta_{21}(x)$ $=\mathbb{1},$ $\zeta_{22}(x)$ $=\sigma_{1}^{z}(x+1).$ (61)
@@
$a\Phi_{1}(x)$ $=\sum_{\alpha}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1),$ (62)
$a\Phi_{2}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (63)
$a\Phi_{3}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (64)
$a^{3/2}\Phi_{4}(x)$ $=6\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{-}(x)\sigma_{2}^{-}(x)\sigma_{3}^{-}(x),$ (65)
where again $\zeta_{\alpha\beta}(x)$ are strings of $Z$ Pauli matrices analogous to the ones of the $SU(2)$ case. Notice that the baryon is a fermionic particle, which implies that the string of Pauli matrices $\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)$ does not cancel out as for bosonic particles. All these operators can be decomposed as a linear combination of unitary operators, and we can implement them with LCU using $O(1)$ ancillary qubits.
@@ V.1 Staggered Fermions in Three Dimensions
where $\mathbf{r}=(x,y,z)$ is a triplet of integers, $\hat{\mathbf{n}}\in\{\hat{\mathbf{x}},\hat{\mathbf{y}},\hat{\mathbf{z}}\}$ is a unit vector along a positive Cartesian axis, and
@@
This Hamiltonian in the continuum limit describes two flavours of quarks, up and down for example, with the same mass $m_{u}=m_{d}=m_{0}$. Identifying the components of the two Dirac spinors $u(\mathbf{r})$ and $d(\mathbf{r})$ with the field $\xi(\mathbf{r})$ is not trivial, and several choices can be made up to $O(a)$ differences. What we propose here is based on [36, 37], where they perform a similar identification in momentum space. First modify $\xi(\mathbf{r})$ by a phase,
@@
$\begin{gathered}f_{1}(\mathbf{r})=\phi(2\mathbf{r}),\quad f_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{z}}),\\ f_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{z}}),\quad f_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}})\\ g_{1}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}),\quad g_{2}(\mathbf{r})=\phi(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\\ g_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{4}(\mathbf{r})=\phi(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}).\end{gathered}$ (70)
@@
$\tilde{u}=\frac{1}{\sqrt{2}}(f+g),\qquad\tilde{d}=\frac{M}{\sqrt{2}}(f-g),$ (71)
$u=\frac{1}{\sqrt{2}}(\tilde{u}-i\tilde{d}),\qquad d=\frac{1}{\sqrt{2}}(-i\tilde{u}+\tilde{d}),$ (72)
@@
$u=\frac{1}{2}\big{[}(\mathbb{1}-iM)f+(\mathbb{1}+iM)g\big{]},\qquad d=\frac{1}{2}\big{[}(M-i\mathbb{1})f-(M+i\mathbb{1})g\big{]}.$ (74)
@@
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}+\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}u(\mathbf{r})-d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}d(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]}\end{gathered}$ (75)
@@
$\begin{gathered}\Delta_{\hat{\mathbf{x}}}q(\mathbf{r})=\begin{pmatrix}q_{1}(\mathbf{r}+\hat{\mathbf{x}})-q_{1}(\mathbf{r})\\ q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{x}})\\ q_{3}(\mathbf{r}+\hat{\mathbf{x}})-q_{3}(\mathbf{r})\\ q_{4}(\mathbf{r})-q_{4}(\mathbf{r}-\hat{\mathbf{x}})\end{pmatrix},\quad\Delta_{\hat{\mathbf{y}}}q(\mathbf{r})=\frac{1}{2}\begin{pmatrix}q_{1}(\mathbf{r})-q_{1}(\mathbf{r}-\hat{\mathbf{y}})+q_{1}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})-q_{1}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{2}(\mathbf{r}+\hat{\mathbf{y}})-q_{2}(\mathbf{r})+q_{2}(\mathbf{r}-\hat{\mathbf{x}})-q_{2}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{y}})+q_{3}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})-q_{3}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{4}(\mathbf{r}-\hat{\mathbf{x}})-q_{4}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})+q_{4}(\mathbf{r}+\hat{\mathbf{y}})-q_{4}(\mathbf{r})\end{pmatrix},\\ \Delta_{\hat{\mathbf{z}}}q(\mathbf{r})=\begin{pmatrix}q_{1}(\mathbf{r}+\hat{\mathbf{z}})-q_{1}(\mathbf{r})\\ q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{4}(\mathbf{r}+\hat{\mathbf{z}})-q_{4}(\mathbf{r})\end{pmatrix},\quad\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}q(\mathbf{r})=\frac{1}{2}\begin{pmatrix}q_{1}(\mathbf{r})-q_{1}(\mathbf{r}-\hat{\mathbf{y}})-q_{1}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})+q_{1}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{2}(\mathbf{r}+\hat{\mathbf{y}})-q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{x}})+q_{2}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{y}})-q_{3}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})+q_{3}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{4}(\mathbf{r}-\hat{\mathbf{x}})-q_{4}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})-q_{4}(\mathbf{r}+\hat{\mathbf{y}})+q_{4}(\mathbf{r})\end{pmatrix}.\end{gathered}$ (77)
@@
$\begin{gathered}f_{1}(\mathbf{r})=\phi(2\mathbf{r}),\quad f_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{z}}),\quad f_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{z}}),\quad f_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}})\\ g_{1}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}),\quad g_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{y}}),\end{gathered}$ (79)
@@
$u=\frac{1}{\sqrt{2}}(f+g),\qquad d=\frac{M}{\sqrt{2}}(f-g).$ (80)

### v2

@@ III Interpolating Operators in One-dimensional Lattice Gauge Theories
In this and the following sections, we aim at showing how to decompose interpolating operators into a linear combination of unitaries on qubit systems, providing some rough estimates on the resources needed and the success probability. We start by considering the simpler case of gauge theories in one space dimension, which is enough to show all the building blocks that are necessary in the physical case of quantum chromodynamics.
The Kogut-Susskind Hamiltonians for lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$ in one space dimension can be compactly written as [70]
@@ III.1 $U(1)$ gauge theory
In this model there are three stable mesons in the continuum limit, with associated interpolating operators [71]
@@
The Dirac spinor $\psi$ for the quark field has two components in one dimension. In the staggered-fermion formulation, the upper component is identified with fermionic modes at even sites, while the lower component is identified with fermionic modes at odd sites. Furthermore, the dimensionless field $\xi$ has to be multiplied by a factor $1/\sqrt{a}$ to match the continuum normalization, $\psi\sim\xi/\sqrt{a}$. The operators in equation (21) translate respectively to
$\hat{\mathcal{O}}_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (22)
$\hat{\mathcal{O}}_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (23)
$\hat{\mathcal{O}}_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (24)
where we have dropped the factor $i$ in $\hat{\mathcal{O}}_{3}$ as it is irrelevant for our purposes.
@@ III.2 $SU(2)$ gauge theory
Analogously to the $U(1)$ theory, the $SU(2)$ theory has mesons [72]
$\hat{\mathcal{O}}_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (36)
$\hat{\mathcal{O}}_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (37)
$\hat{\mathcal{O}}_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]}.$ (38)
@@
$\hat{\mathcal{O}}_{4}(x)=\frac{1}{a}\epsilon_{\alpha\beta}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x),$ (39)
@@ III.3 $SU(3)$ gauge theory
The $SU(3)$ case is a more complicated version of the $SU(2)$ one. For the details see [70]. The basis states are labeled by eight quantum numbers
@@
$\hat{\mathcal{O}}_{4}(x)=\frac{1}{a^{3/2}}\epsilon_{\alpha\beta\gamma}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x)\xi_{\gamma}^{\dagger}(x).$ (47)
@@ IV Interpolating Operators on Qubits
$\left.\begin{aligned} \xi_{1}(x)&\rightarrow\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{2}^{z}(x-1)\sigma_{1}^{+}(x)\\ \xi_{2}(x)&\rightarrow\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{2}^{z}(x-1)\sigma_{1}^{z}(x)\sigma_{2}^{+}(x)\end{aligned}\right\}$ for $SU(2)$, (49)
$\left.\begin{aligned} \xi_{1}(x)&\rightarrow\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{+}(x)\\ \xi_{2}(x)&\rightarrow\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{z}(x)\sigma_{2}^{+}(x)\\ \xi_{3}(x)&\rightarrow\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{z}(x)\sigma_{2}^{z}(x)\sigma_{3}^{+}(x)\end{aligned}\right\}$ for $SU(3)$, (50)
where $\sigma^{\pm}=(\sigma^{x}\pm i\sigma^{y})/2$.
On the other hand, the local Hilbert space of the gauge field needs to be truncated. For $U(1)$ this is quite simple: if we use $k$ qubits for a single link, then we impose the cutoff $\Lambda=2^{k-1}$ and we encode the states $|-\Lambda\rangle,|-\Lambda+1\rangle,\dots,|\Lambda-1\rangle$ in the computational basis. The operator $U$ needs to be adapted to the truncation. A common choice is to take
@@
In this way $U$ is a permutation operator and can be implemented on a qubit system using $O(k^{3})$ gates and no ancillary qubit, or using $O(k^{2})$ gates and k1k-1 ancillary qubits. For $SU(2)$ and $SU(3)$ the situation is more complicated but similar in essence. We refer the reader to [70] for the details. After imposing a cutoff $\Lambda$, the permutation operators $M_{\alpha}^{\text{L}/\text{R}},\,J^{\pm}$ of $SU(2)$, and $M_{\alpha}^{\text{L}/\text{R}},\,P^{\pm},\,Q^{\pm},\,T_{\text{L}/\text{R}}^{\pm}$ of $SU(3)$, take the same form as $U$ in equation (51).
The diagonal operators $c_{\alpha\beta}^{\pm}$ in (31) and the ones in (43), like for instance $\hat{C}_{11}^{\text{L}(a)}\hat{C}_{11}^{\text{R}(a)}\hat{N}_{1}$, are not unitary. However, their entries lie between 1-1 and $1$, and are efficiently computable classically. We can implement them in the following way: suppose that the function $f(x)$, where $x$ can take $M$ values encoded in the computational basis, is efficiently computable with classical techniques, and take an ancillary register of $\eta=O(\log_{2}M)$ qubits initialized in the state
@@
We can now write the interpolating operators in the qubit system, together with the corresponding quantity $C$ defined in equation (8). For $U(1)$ they take the form
$\hat{\mathcal{O}}_{1}(x)$ $=\frac{1}{a}\big{[}\sigma^{+}(x)\sigma^{-}(x)-\sigma^{+}(x+1)\sigma^{-}(x+1)\big{]},$ $\quad C_{1}$ $=\frac{2}{a},$ (55)
$\hat{\mathcal{O}}_{2}(x)$ $=\frac{1}{a}\big{[}\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)+\sigma^{-}(x+1)U(x)\sigma^{+}(x)\big{]},$ $\quad C_{2}$ $=\frac{2}{a},$ (56)
$\hat{\mathcal{O}}_{3}(x)$ $=\frac{1}{a}\big{[}\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)-\sigma^{-}(x+1)U(x)\sigma^{+}(x)\big{]},$ $\quad C_{3}$ $=\frac{2}{a}.$ (57)
@@
$\hat{\mathcal{O}}_{1}(x)$ $=\frac{1}{a}\sum_{\alpha}\big{[}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1)\big{]},$ $\quad C_{1}$ $=\frac{4}{a},$ (58)
$\hat{\mathcal{O}}_{2}(x)$ $=\frac{1}{a}\sum_{\alpha,\beta}\big{[}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x)\big{]},$ $\quad C_{2}$ $=\frac{16}{a},$ (59)
$\hat{\mathcal{O}}_{3}(x)$ $=\frac{1}{a}\sum_{\alpha,\beta}\big{[}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x)\big{]},$ $\quad C_{3}$ $=\frac{16}{a},$ (60)
$\hat{\mathcal{O}}_{4}(x)$ $=\frac{2}{a}\sigma_{1}^{-}(x)\sigma_{2}^{-}(x),$ $\quad C_{4}$ $=\frac{2}{a},$ (61)
@@
$\zeta_{21}(x)$ $=\mathbb{1},$ $\zeta_{22}(x)$ $=\sigma_{1}^{z}(x+1),$ (63)
and the $U_{\alpha\beta}$ are decomposed as in equation (31).
@@
$\hat{\mathcal{O}}_{1}(x)$ $=\frac{1}{a}\sum_{\alpha}\big{[}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1)\big{]},$ $\quad C_{1}$ $=\frac{6}{a},$ (64)
$\hat{\mathcal{O}}_{2}(x)$ $=\frac{1}{a}\sum_{\alpha,\beta}\big{[}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x)\big{]},$ $\quad C_{2}$ $=\frac{130}{a},$ (65)
$\hat{\mathcal{O}}_{3}(x)$ $=\frac{1}{a}\sum_{\alpha,\beta}\big{[}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x)\big{]},$ $\quad C_{3}$ $=\frac{130}{a},$ (66)
$\hat{\mathcal{O}}_{4}(x)$ $=\frac{6}{a^{3/2}}\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{-}(x)\sigma_{2}^{-}(x)\sigma_{3}^{-}(x),$ $\quad C_{4}$ $=\frac{6}{a^{3/2}},$ (67)
where again $\zeta_{\alpha\beta}(x)$ are strings of $Z$ Pauli matrices analogous to the ones of the $SU(2)$ case, and the $U_{\alpha\beta}$ are decomposed as in equation (43). Notice that the baryon is a fermionic particle, which implies that the string of Pauli matrices $\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)$ does not cancel out as for bosonic particles.
We have thus provided explicit linear expansions of the interpolating operators in terms of unitary operators on qubit systems. The number of terms varies considerably from theory to theory and from particle to particle, but in all cases it is a constant. In the cases of $SU(2)$ and $SU(3)$, the quantities $C_{2}$ and $C_{3}$ are quite large because of the appearance of the operators $U_{\alpha\beta}$ in the expansion. This feature is specific of the Kogut-Susskind formulation. The number of gates required for each term in the expansion grows at most polynomially with the number of qubits dedicated to a lattice site.
@@ V Quantum Chromodynamics
We focus now on two-flavoured quantum chromodynamics in three dimensions. The first thing we need is to identify the continuum degrees of freedom with the lattice degrees of freedom. This is what we do in subsection V.1. Then, in subsection V.2 we discuss the wave-packet creation of pions and nucleons.
@@ V.1 Staggered fermions in three dimensions
[…]
@@
This Hamiltonian in the continuum limit describes two flavours of quarks, up and down for example, with the same mass $m_{u}=m_{d}=m_{0}$. Identifying the components of the two Dirac spinors $u(\mathbf{r})$ and $d(\mathbf{r})$ with the field $\xi(\mathbf{r})$ is not trivial, and several choices can be made up to $O(a)$ differences. We show here two possibilities based on [73, 74], where they perform a similar identification in momentum space.
First modify $\xi(\mathbf{r})$ by a phase,
@@
[…]
@@
$\tilde{u}=\frac{1}{\sqrt{2}a^{3/2}}(f+g),\qquad\tilde{d}=\frac{M}{\sqrt{2}a^{3/2}}(f-g),$ (73)
[…]
@@
$u=\frac{1}{2a^{3/2}}\big{[}(\mathbb{1}-iM)f+(\mathbb{1}+iM)g\big{]},\qquad d=\frac{1}{2a^{3/2}}\big{[}(M-i\mathbb{1})f-(M+i\mathbb{1})g\big{]}.$ (76)
@@
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}a^{3}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}+\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}u(\mathbf{r})-d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}d(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]}\end{gathered}$ (77)
@@
[…]
@@
[…]
@@
$u=\frac{1}{\sqrt{2}a^{3/2}}(f+g),\qquad d=\frac{M}{\sqrt{2}a^{3/2}}(f-g).$ (82)